OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4.

OpenStudy (whpalmer4):

\[3(x+1) = -2(x-1) -4\]Start by distributing everything: \[3*x + 3*1 = -2*x -2*(-1) - 4\]\[3x+3=-2x+2-4\]\[3x+3=-2x-2\]Now add \(2x\) to both sides and subtract \(3\) from both sides. You'll be left with an equation that you can solve for \(x\) by dividing by the number in front of the \(x\) (the coefficient).